Given the demand for a product as Qd = 100 - 5P and the supply is given as Qs = -20 + 3p. You are told equilibrium is obtained at the point where Qd = Qs. The equilibrium quantity for the product is

1
Expert's answer
2021-06-15T12:32:19-0400

"Q_d=Q_s\\100-5P=-20+3P\\100+20=3P+5P\\120=8P\\P=15"


"Q_d=100-5(15)=25\\Q_s=-20+3(15)=25"


Therefore the quantity of the product is 25.